Define the following with two examples
Zeros function
Even function
Odd function
Heaviside function
Greatest integer function
Absolute function
Identity function
Constant function
Periodic function
1
Expert's answer
2019-08-19T13:38:46-0400

1) Zero function: This is a function equal to zero on every point of its domain ("y=0, y=0(x+1)" );

2) Even function: For every x of this type of function "f(-x)=f(x)" ("y=x^2, y=cos(x))"

3) Odd function: For every x of this type of function "f(-x)=-f(x)" ("y=x^3, y=sin(x)")

4) Heaviside function: This is a unit step function "H(x):=\frac{d}{dx}max(x,0)" ("H(x)=\frac{1}{2}+\frac{1}{2}sgn(x), H(x)=1_{[0, infinity)}(x)")

5) Greatest integer function: The Greatest integer function is defined as the largest integer which is less than or equal to x ("y=[x], y=\frac{1}{2}[x]")

6) Absolute number: This is a positive number for positive x and the opposite number for a negative x ("y=\mid{x}\mid, y=\mid{x+1}\mid)"

7) Identity function: This is actually the same value that was used as an argument ("y=x")

8) Constant function: This is the function with the same value for every argument x ("y=3, y=3.14")

9) Periodic function: For every x of this function "f(x)=f(x+a)", where a is a non-zero real number or period ("y=sin(x), y=cos(x)")